### v1

@@ II.3 Temperature Derivatives
 $\left.\mbox{ }-[-2{\cal H}\dot{W}_{\rm i}+2\dot{W}_{\rm r}\dot{W}_{\rm i}+\ddot{W}_{\rm i}]\tan W_{\rm i}\right\rangle_{W}.$
Note that the heat capacity is $C_{V}=\partial\tilde{E}/\partial T=-k_{\mathrm{B}}\beta^{2}\partial\tilde{E}/\partial\beta$.
@@ III.2 Results
At $T^{*}=0.5$ the effective energy is $\beta\tilde{E}/N=-10.15(1)$, and the heat capacity is $C_{V}/Nk_{\rm B}=39.6(19)$. At $T^{*}=0.45$ these are $\beta\tilde{E}/N=-19.49(2)$ and $C_{V}/Nk_{\rm B}=59.7(23)$.
@@
Real 4He does not become solid at saturation pressure down to absolute zero, but it does solidify at elevated pressures. One might speculate that the present solid-like state results from a combination of two effects. The first is the approximate nature of the Lennard-Jones pair potential together with the termination of the temperature expansion for the commutation function at $n_{W}^{\rm max}=3$. The second is the Laplace pressure acting on the interior of the nano-droplet.

### v2

@@ II.3 Temperature Derivatives
 $\left.\mbox{ }-[-2{\cal H}\dot{W}_{\rm i}+2\dot{W}_{\rm r}\dot{W}_{\rm i}+\ddot{W}_{\rm i}]\tan W_{\rm i}\right$.
 $\left.\mbox{ }-\dot{W}_{\rm i}^{2}\sec^{2}W_{\rm i}\right\rangle_{W}.$
(The final term was missing in version 1.) Note that the heat capacity is $C_{V}=\partial\tilde{E}/\partial T=-k_{\mathrm{B}}\beta^{2}\partial\tilde{E}/\partial\beta$.
@@ III.2 Results
For $k_{\rm B}T/\varepsilon=0.50$ and $\rho_{\rm tot}\sigma^{3}=0.26$ (droplet, central density 0.290(9), edge oscillations), gives $\beta\overline{E}/N=-12.41(4)$ and $\beta^{2}\partial\overline{E}/N\partial\beta=34(2)$. For $k_{\rm B}T/\varepsilon=0.49$, $\beta\overline{E}/N=-13.41(4)$ and $\beta^{2}\partial\overline{E}/N\partial\beta=37(1)$. (These results have been recalculated from version 1.)
The numerical difference is $\Delta E/Nk_{\rm B}\Delta T=37.06(3)$. That this agrees with the directly simulated heat capacity is a very sensitive test of the mathematical expressions for the energy and the heat capacity, and of their computer implementation. It also tests the Metropolis Monte Carlo algorithm and its implementation. The author can confirm that it takes many rounds of debugging to get quantitative agreement between the two.
The energy and heat capacity are sensitive to the saturated droplet and vary with the total density. For example, at $k_{\rm B}T/\varepsilon=0.50$, an overall density $\rho_{\rm tot}\sigma^{3}=0.15$. forms a smooth droplet with central density 0.28 and gives $\beta\overline{E}/N=-10.37(1)$. An overall density $\rho_{\rm tot}\sigma^{3}=0.03$ forms a smooth droplet with central density 0.28 and gives $\beta\overline{E}/N=-8.6(1)$. In these the energy of all atoms in the system are included.
@@
Real 4He does not become solid at saturation pressure down to absolute zero, but it does solidify at elevated pressures. One might speculate that the present solid-like state results from a combination of three effects. The first is the approximate nature of the Lennard-Jones pair potential together with the termination of the temperature expansion for the commutation function at $n_{W}^{\rm max}=3$. The second is the Laplace pressure acting on the interior of the nano-droplet. The third is the periodic boundary conditions.